OpenStudy (anonymous):

Can some one please explain Graphing Linear Equations to me? No question specificaly, just everything you know about it? Please?

OpenStudy (anonymous):

Now I have a specific question! Are all points on a line are solutions to the equation of the line?

OpenStudy (anonymous):

^^^^^^^^ True/false question

OpenStudy (anonymous):

read the attachment please!

OpenStudy (whpalmer4):

Yes, all points on a line are solutions to the equation of the line. The bottom line with graphing linear equations is that the graphs are always straight lines.

OpenStudy (whpalmer4):

The name might be a hint in this regard :-)

OpenStudy (anonymous):

Thanks!!!! Can you help with the other problems on the attachment?

OpenStudy (compassionate):

I can help you.

OpenStudy (compassionate):

1 For which equation is (4, 3) a solution? y = 2x – 1 y = 3x – 4 y = 2x – 5 y = x + 3 Just plug in (4, 3) to each equation. If the statement is true, then it's a solution. E.G: y = 2x - 1 3 = 2(4) - 1 3 = 8 - 1 3= 7, this is not a true statement and thus false.

OpenStudy (anonymous):

Thanks!

OpenStudy (compassionate):

DO you know how to solve the rest?

OpenStudy (anonymous):

I would like help with 2, but I'm pretty stable with the rest.

OpenStudy (anonymous):

Nevermind, i figured out the answer. Thanks anyway!

OpenStudy (compassionate):

Good luck.

OpenStudy (anonymous):

thanks.